# MATHWELL: Generating Educational Math Word Problems at Scale

### 28 Feb 2024 v2

@@ 1 Introduction
Math word problems include natural language questions paired with numerical solutions, as illustrated in Figure 1. To be effective educational tools for grade-school students, the problems must be solvable, appropriate, and have accurate solutions. While most public word problem resources are manually curated and, therefore, quite limited (e.g., K5Learning ; TPT ; VDOE ), there have been recent efforts to generate math word problems using natural language processing Jiao et al. (2023); Koncel-Kedziorski et al. (2016); Niyarepola et al. (2022); Norberg et al. (2023); Qin et al. (2023); Wang et al. (2021); Zhou and Huang (2019); Zhou et al. (2023); Zong and Krishnamachari (2023). Closest to our work, some recent efforts explore math word problem generation using Large Language Models (LLMs) Niyarepola et al. (2022); Zong and Krishnamachari (2023). However, to the best of our knowledge, existing methods have one or both of two important limitations. First, many only generate questions (e.g., Niyarepola et al. 2022; Zong and Krishnamachari 2023). To be educationally useful, students and teachers must also have access to answers. Further, it is hard to verify if questions are solvable and appropriate without answers. Second, all existing methods are context-dependent, simply rephrasing or creating new output based on input problems or equations. In practice, curating and choosing specific word problems or equations beforehand is laborious and difficult Roche (2013).
To address the aforementioned limitations, we propose context-free educational math word problem generation. Successful solutions must return questions and answers together and require no hand-picked equations or exemplar word problems. For context-free word problem generators to be effective educational tools, we propose three criteria to evaluate their outputs. 1) Each problem must be solvable. 2) Each question’s corresponding answer must be accurate. 3) Each problem must be appropriate, where the math topic should be familiar to a student and the question’s context should be appropriate for a young learner.
@@
We use MATHWELL to generate Synthetic Grade School Math (SGSM), a dataset of 20,490 word problems. This is the largest public set of English math word problems with PoT rationales. We acquire annotations from domain experts for 3,484 model samples during and after training to score the quality of our data, resulting in a SGSM subset with 2,093 samples with gold labels for solvability, accuracy, and appropriateness. We find that MATHWELL is effective at generating custom math word problems without context: 74% of its output problems with executable code are simultaneously solvable, accurate, and appropriate and nearly all successfully include a specified topic. These outputs are similar to human-written math QA datasets. MATHWELL has a nearly 20% higher share of outputs with executable code that meet our three evaluation criteria than alternative LLMs. Further, we find that SGSM has a more age appropriate reading level than other datasets. We also show that our human annotations can be used to train text classifiers to automatically score MATHWELL outputs.
@@ 2 Related Work
There are several existing datasets for training LLMs to answer grade school math questions in English, including GSM8K Cobbe et al. (2021), NumGLUE Mishra et al. (2022b), GSM-Hard Gao et al. (2023), ASDIV Miao et al. (2020) and SVAMP Patel et al. (2021). However, because these datasets were created to train LLMs to answer, rather than generate, grade school math problems, they do not possess the characteristics we consider critical to training LLMs to generate solvable, accurate, and appropriate question/answer pairs. Specifically, while some datasets include PoT solutions (GSM-Hard, MathInstruct GSM8K, and NumGLUE), only GSM-Hard contains Python function solutions, which we find increases the likelihood that a question/answer pair contains executable code. Because GSM-Hard contains questions with inflated number values, though, it is not an appropriate dataset for training a question generator for K-8 students. Due to the limitations of existing datasets, we create and release SGSM, a SGSM subset verified by human annotation that contains Python function solutions and appropriate number ranges for K-8 students.
@@
Other work explores automatically generating math word problems but takes a more limited approach than the present study. The closest related work is Zong and Krishnamachari (2023), who assess GPT-3’s ability to generate math word problems. Their method is not context-free, however, as they use a reference problem to guide generation. In addition, they only generate word problems, not solutions. Niyarepola et al. (2022) train LLM word problem generators, but they guide generation by providing the beginning of a word problem and do not generate solutions. Most other existing works use LLMs or deep neural networks to generate word problems based on pre-specified equations, provide additional questions based on reference word problems, or re-write existing word problems Jiao et al. (2023); Koncel-Kedziorski et al. (2016); Norberg et al. (2023); Qin et al. (2023); Wang et al. (2021); Zhou and Huang (2019); Zhou et al. (2023), each of which are more limited than context-free generation by restricting the range of possible outputs. These approaches also require additional input from users, which may not be feasible for teachers or students who wish to create customized word problems. As a result, we do not compare MATHWELL against these models/methods. Our study is the first to evaluate the capability of leading open-source LLMs for the task of math word problem generation.
@@ 3 Methods
There are three critical criteria for model-generated question/answer pairs: 1) Solvability indicates questions are possible to solve and have one correct answer Jiao et al. (2023); Koncel-Kedziorski et al. (2016); Qin et al. (2023). 2) Accuracy measures if the generated answer is correct. 3) Appropriateness indicates that questions make logical sense, do not have grammatical errors or conflicting information, and are appropriate for K-8 students mathematically and contextually. A secondary characteristic we consider for model comparisons is topic specificity, or if the question incorporates the topic specified in the prompt.
@@ 3.1 Training MATHWELL
Inspired by recent works that iteratively finetune LLMs Guo et al. (2024); Wang et al. (2024), including OpenAI’s FeedME approach Zheng et al. (2023), we next aim to evaluate and improve how well MATHWELL meets our three main criteria: solvability, accuracy, and appropriateness. However, these properties are unannotated in prior works. Therefore, we few-shot prompt the finetuned model and generate 3,234 problems. We then acquire annotations for solvability, accuracy, and appropriateness from volunteers with previous K-12 teaching experience/training. Further details about the annotation process and annotator agreement are in Appendix F. These annotations identified 1,906 problems that are simultaneously solvable, accurate, and appropriate, or meet all criteria (MaC). We then finetune MATHWELL on these MaC outputs for 1,250 more steps.
@@ 4 Experiments
We evaluate MATHWELL’s outputs using both human evaluations with domain experts and automatic evaluations. We compare 250 randomly-selected MATHWELL outputs to those of LLEMMA (34B) Azerbayev et al. (2023), MAmmoTH (70B) Yue et al. (2023), and Llama-2 (70B) Touvron et al. (2023), which are recent and popular. LLEMMA and MAmmoTH represent two different approaches to training math LLMs, as MAmmoTH is trained as a QA model Yue et al. (2023), while LLEMMA is trained to be more general using a large corpus of mathematical content Azerbayev et al. (2023). Llama-2 serves as a baseline to ensure task-specific finetuning increases performance. Each model is prompted using example question/answer pairs from SGSM and asked to create a question/answer pair on a random topic K-8 students are interested in. Our 250 evaluation sample for each model is roughly five times larger than the human evaluation samples used in other math word problem generator studies Jiao et al. (2023); Koncel-Kedziorski et al. (2016); Niyarepola et al. (2022); Zhou and Huang (2019); Zhou et al. (2023); Zong and Krishnamachari (2023).
@@ E.4.3 Appropriateness
Appendix F Annotation Process
F.1 Annotators
@@
F.2 Inter-Annotator Agreement
@@
F.3 Handling Annotator Disagreement
@@
F.4 Validating Final Evaluation Labels

### 01 Mar 2024 v3

@@ 1 Introduction
Math word problems include natural language questions paired with numerical solutions, as illustrated in Figure 1. To be effective educational tools for grade-school students, the problems must be solvable, appropriate, and have accurate solutions. While most public word problem resources are manually curated and, therefore, quite limited (e.g., K5Learning ; TPT ; VDOE ), there have been recent efforts to generate math word problems using natural language processing Jiao et al. (2023); Koncel-Kedziorski et al. (2016); Niyarepola et al. (2022); Norberg et al. (2023); Qin et al. (2023, 2024); Wang et al. (2021); Zhou and Huang (2019); Zhou et al. (2023); Zong and Krishnamachari (2023). Closest to our work, some recent efforts explore math word problem generation using Large Language Models (LLMs) Niyarepola et al. (2022); Zong and Krishnamachari (2023). However, to the best of our knowledge, existing methods have one or both of two important limitations. First, many only generate questions (e.g., Niyarepola et al. 2022; Zong and Krishnamachari 2023). To be educationally useful, students and teachers must also have access to answers. Further, it is hard to verify if questions are solvable and appropriate without answers. Second, all existing methods are context-dependent, simply rephrasing or creating new output based on input problems or equations. In practice, curating and choosing specific word problems or equations beforehand is laborious and difficult Roche (2013).
To address the aforementioned limitations, we propose context-free educational math word problem generation. Successful solutions must return questions and answers together and require no hand-picked equations or exemplar word problems. For context-free word problem generators to be effective educational tools, we propose three criteria to evaluate their outputs. 1) Each problem must be solvable. 2) Each question’s corresponding answer must be accurate. 3) Each problem must be appropriate, where the question/answer should make logical sense, the math topic should be familiar to a student, and the question’s context should be appropriate for a young learner.
@@
We use MATHWELL to generate Synthetic Grade School Math (SGSM), a dataset of 20,490 word problems. This is the largest public set of English math word problems with PoT rationales. We acquire annotations from domain experts for 3,484 model samples during and after training to score the quality of our data, resulting in a SGSM subset with 2,093 samples with gold labels for solvability, accuracy, and appropriateness. We find that MATHWELL is effective at generating custom math word problems without context: 74% of its output problems with executable code are simultaneously solvable, accurate, and appropriate and nearly all successfully include a specified topic. These outputs are similar to popular human-written math QA datasets. MATHWELL has a nearly 20% higher share of outputs with executable code that meet our three evaluation criteria than alternative LLMs. Further, we find that SGSM has a more age appropriate reading level than other datasets. We also show that our human annotations can be used to train text classifiers to automatically score MATHWELL outputs.
@@ 2 Related Work
There are several popular datasets for training LLMs to answer grade school math questions in English, including GSM8K Cobbe et al. (2021), NumGLUE Mishra et al. (2022b), GSM-Hard Gao et al. (2023), ASDIV Miao et al. (2020) and SVAMP Patel et al. (2021). Recent work has also developed new grade school math datasets or different rationales for existing datasets, including those that are human written Kim et al. (2023) and synthetic Chern et al. (2023); Mitra et al. (2024); Shi et al. (2023); Yu et al. (2023); Yue et al. (2023). However, because existing datasets were created to train LLMs to answer, rather than generate, grade school math problems, they do not possess the characteristics we consider critical to training LLMs to generate solvable, accurate, and appropriate question/answer pairs. Specifically, while some datasets include PoT solutions (e.g., GSM-Hard, MathInstruct GSM8K, and NumGLUE), only GSM-Hard contains Python function solutions, which we find increases the likelihood that a question/answer pair contains executable code. Because GSM-Hard contains questions with inflated number values, though, it is not appropriate for training a question generator for K-8 students. Due to the limitations of existing datasets, we create and release SGSM, a SGSM subset verified by human annotation that contains Python function solutions and appropriate number ranges for K-8 students.
@@
Other work explores automatically generating math word problems but takes a more limited approach than the present study. The closest related work is Zong and Krishnamachari (2023), who assess GPT-3’s ability to generate math word problems. Their method is not context-free, however, as they use a reference problem to guide generation. In addition, they only generate word problems, not solutions. Niyarepola et al. (2022) train LLM word problem generators, but they guide generation by providing the beginning of a word problem and do not generate solutions. Most other existing works use LLMs or deep neural networks to generate word problems based on pre-specified equations, provide additional questions based on reference word problems, or re-write existing word problems Jiao et al. (2023); Koncel-Kedziorski et al. (2016); Norberg et al. (2023); Qin et al. (2023, 2024); Wang et al. (2021); Zhou and Huang (2019); Zhou et al. (2023), each of which are more limited than context-free generation by restricting the range of possible outputs. These approaches also require additional input from users, which may not be feasible for teachers or students who wish to create customized word problems. As a result, we do not compare MATHWELL against these models/methods. Our study is the first to evaluate the capability of leading open-source LLMs for the task of math word problem generation.
@@ 3 Methods
There are three critical criteria for model-generated question/answer pairs: 1) Solvability indicates questions are possible to solve and have one correct answer Jiao et al. (2023); Koncel-Kedziorski et al. (2016); Qin et al. (2023). 2) Accuracy measures if the generated answer is correct. 3) Appropriateness indicates that questions/answers make logical sense, do not have grammatical errors or conflicting information, and are appropriate for K-8 students mathematically and contextually. A secondary characteristic we consider for model comparisons is topic specificity, or if the question incorporates the topic specified in the prompt.
@@ 3.1 Training MATHWELL
Inspired by recent works that iteratively finetune LLMs Guo et al. (2024); Wang et al. (2024), including OpenAI’s FeedME approach Zheng et al. (2023), we next aim to evaluate and improve how well MATHWELL meets our three main criteria: solvability, accuracy, and appropriateness. However, these properties are unannotated in prior works. Therefore, we few-shot prompt the finetuned model and generate 3,234 problems. We then acquire annotations for solvability, accuracy, and appropriateness from volunteers with previous K-12 teaching experience/training. Further details about the annotation process and annotator agreement are in Appendix G. These annotations identified 1,906 problems that are simultaneously solvable, accurate, and appropriate, or meet all criteria (MaC). We then finetune MATHWELL on these MaC outputs for 1,250 more steps.
@@ 4 Experiments
We evaluate MATHWELL using both human evaluations with domain experts and automatic evaluations. We compare 250 randomly-selected MATHWELL outputs to those of LLEMMA (34B) Azerbayev et al. (2023), MAmmoTH (70B) Yue et al. (2023), and Llama-2 (70B) Touvron et al. (2023), which are recent and popular. LLEMMA and MAmmoTH represent two different approaches to training math LLMs, as MAmmoTH is trained as a QA model Yue et al. (2023), while LLEMMA is trained more generally using a corpus of math content Azerbayev et al. (2023). Llama-2 serves as a baseline to ensure task-specific finetuning improves performance. Each model is prompted using examples from SGSM and asked to create a question/answer pair on a random topic K-8 students are interested in. Our 250 evaluation sample for each model is roughly 2.5-5 times larger than the human evaluation samples used in other word problem generator studies Jiao et al. (2023); Koncel-Kedziorski et al. (2016); Niyarepola et al. (2022); Qin et al. (2024); Zhou and Huang (2019); Zhou et al. (2023); Zong and Krishnamachari (2023).
@@ E.4.3 Appropriateness
Appendix F Early MATHWELL Experimentation
Appendix G Annotation Process
G.1 Annotators
@@
G.2 Inter-Annotator Agreement
@@
G.3 Handling Annotator Disagreement
@@
G.4 Validating Final Evaluation Labels